### v1

Abstract
Recent years have witnessed great progress in image restoration thanks to the advancements in modern deep neural networks e.g. Convolutional Neural Network and Transformer. However, existing restoration backbones are usually limited due to the inherent local reductive bias or quadratic computational complexity. Recently, Selective Structured State Space Model e.g., Mamba, has shown great potential for long-range dependencies modeling with linear complexity, but it is still under-explored in low-level computer vision. In this work, we introduce a simple but strong benchmark model, named MambaIR, for image restoration. In detail, we propose the Residual State Space Block as the core component, which employs convolution and channel attention to enhance the capabilities of the vanilla Mamba. In this way, our MambaIR takes advantage of local patch recurrence prior as well as channel interaction to produce restoration-specific feature representation. Extensive experiments demonstrate the superiority of our method, for example, MambaIR outperforms Transformer-based baseline SwinIR by up to 0.36dB, using similar computational cost but with a global receptive field. Code is available at https://github.com/csguoh/MambaIR.
Keywords:
Image restoration State space model Mamba
1 Introduction
Image restoration, aiming to reconstruct a high-quality image from a given low-quality input, is a long-standing problem in computer vision and further has a wide range of sub-problems such as super-resolution, image denoising, etc. Since several pioneering works [64, 65, 13], Convolutional Neural Networks (CNNs) have gained great success than traditional model-based method [21, 10], and almost dominate this field in the past few years [30, 56, 35, 67, 11, 63].
Figure 1: The Effective Receptive Field (ERF) visualization [38] for EDSR [35], RCAN [68], SwinIR [34], HAT [7], and the proposed MambaIR. A larger ERF is indicated by a more extensively distributed dark area. Only the proposed MambaIR achieves a significant global effective receptive field.
Despite the promising results, these restoration backbones usually suffer from the inherent problem of CNNs. First, the weight-sharing strategy in CNN makes the model remain static for varying inputs, preventing robustness on unseen real-world low-quality images. Second, the receptive field of one convolution layer is constrained by the kernel size, which leads to strong local reductive bias. Although stacking multiple convolutional layers can theoretically enlarge the receptive field, however as shown in Fig. 1(a) the effective receptive field [38] of CNN-based restoration methods is still limited.
Inspired by the success in natural language processing, Transformer [54] has proven superior performance than CNNs in several vision tasks due to weak inductive bias as well as global interaction modeling [4, 14, 37]. However, taming Transformer for image restoration poses a significant challenge since the self-attention mechanism scales quadratically with the number of tokens. Considering that image restoration methods usually set the number of tokens to the image resolution, therefore directly using the attention mechanism will incur a huge computational overhead. Although some works make efforts to divide input into small patches [6] or employ shifted window attention [34, 7], these strategies harm the long-term dependency [7] as shown in Fig. 1(b), and do not essentially address the quadratic complexity problem.
Recently, structured state-space sequence models (S4), have emerged as an efficient and effective backbone for constructing deep networks [18, 52, 15, 43, 17]. In particular, the improved S4 with selective mechanism and efficient hardware design, also called Mamba, has been proven to surpass Transformer on tasks requiring long-term dependency modeling such as natural language processing [17]. More recently, some variants of Mamba have also presented promising results in computer vision [72, 36, 40], such as image classification, biomedical image segmentation, etc. However, the power of Mamba in low-level vision is still under-explored, which motivates us to explore the potential of Mamba to enhance the long-range modeling ability for image restoration networks.
In this work, we propose a simple but effective benchmark restoration model, namely MambaIR, to adapt Mamba for image restoration. Following  [34, 7, 62], our MambaIR consists of three stages: shallow feature extraction, deep feature extraction, and high-quality image reconstruction. The shallow feature extraction employs a simple convolution layer to extract the shallow feature. Then the deep feature extraction is performed with several stacked Residual State Space Blocks (RSSBs). As the core component of MambaIR, the RSSB is designed to enhance the capabilities of the original Mamba with image restoration priors. Specifically, to tackle the problem that the original Mamba treats each token equally, ignoring the patch recurrence occurring in the neighborhood, we add additional convolution layers to give local pixels more focus. We further introduce channel attention layer [22] to supplement channel interaction which is beneficial for image restoration but missed in Mamba. In addition, we use the skip connection within each block to offer shortcuts for feature aggregation. Finally, the high-quality image reconstruction stage fuses both shallow and deep features to restore the high-quality output image.
Our main contributions can be summarized as follows:
  • We are the first to work to adapt state space models for low-level image restoration via extensive experiments to formulate MambaIR, which acts as a simple but effective alternative for CNNs and Transformers.
  • We propose the Residue State Space Block (RSSB) which can boost the power of the original Mamba with local spatial prior and channel interaction.
  • Extensive experiments on various tasks demonstrate our MambaIR outperform strong Transformer-based baselines to provide a strong and promising backbone model solution for image restoration.
2 Related Work
2.1 Image Restoration
Image restoration has been significantly advanced since the introduction of deep learning by several pioneering works, such as SRCNN [13] for image super-resolution, DnCNN [64] for image denoising and ARCNN [12] for JPEG compression artifact reduction, etc. Early attempts usually elaborate CNNs with techniques such as residual connection [27, 63, 5], dense connection [56, 69] and others [29, 58, 16, 68, 11] to improve model representation ability. Despite the success, CNN-based restoration methods typically face challenges in effectively modeling global dependencies as well as static weight for varying inputs. As Transformers have been proven to be a strong competitor for CNNs in multiple computer vision tasks [4, 14, 37], using Transformer for image restoration appears promising. However, it still faces specific challenges from the quadratic computational complexity of the self-attention mechanism [54]. To address this, IPT [6] divides one image into several small patches and processes each patch independently with self-attention. SwinIR [34] further introduces shifted window attention to improve the performance. In addition, promising progress continues to be made in designing efficient and effective attention mechanisms for restoration [62, 7, 31, 9, 59, 8]. Nonetheless, efficient attention design usually comes at the expense of global effective receptive fields, and the issue of quadratic computational complexity is not essentially addressed.
2.2 State Space Models
State Space Models (SSMs) [19, 18, 52], stemming from classics control theory [26], are recently introduced to deep learning as a competitive backbone for state space transforming. The promising property of linearly scaling with sequence length in long-range dependency modeling has attracted great interest from searchers. For example, the Structured State-Space Sequence model (S4) [18] is a pioneer work for the deep state-space model in modeling the long-range dependency. Later, S5 layer [52] is proposed based on S4 and introduces MIMO SSM and efficient parallel scan. Moreover, H3 [15] achieves promising results that nearly fill the performance gap between SSMs and Transformers in natural language.  [43] further improve S4 with gating units to obtain the Gated State Space layer to boost the capability. More recently, Mamba [17], a data-dependent SSM with selective mechanism and efficient hardware design, outperforms Transformers on natural language and enjoys linear scaling with input length. And there are also pioneering works that adopt Mamba to vision tasks such as image classification [36, 72], video understanding [55], biomedical image segmentation [40] and others [25, 47]. In this work, we explore the potential of Mamba to image restoration with restoration-specific prior to sever as a simple but effective baseline for future work.
3 Methodology
This work aims to introduce the recent advanced state space model (SSM), i.e., Mamba, to low-level image restoration. In this section, we begin with a description of the preliminaries of the SSM. Then we give an overview of our simple but effective benchmark model MambaIR, followed by a detailed description of the proposed Residule State-Space Block (RSSB). Finally, we provide a comprehensive discussion of the difference between our MambaIR and previous methods.
3.1 Preliminaries
The recent advancements of the class of structured state-space sequence models (S4) are largely inspired by the continuous linear time-invariant (LTI) systems, which maps a 1-dimensional function or sequence $x(t)\in\mathbb{R}\rightarrow y(t)\in\mathbb{R}$ through an implicit latent state $h(t)\in\mathbb{R}^{N}$. Formally, this system can be formulated as a linear ordinary differential equation (ODE):
$h^{\prime}(t)$ $={\rm\textbf{A}}h(t)+{\rm\textbf{B}}x(t),$ (1)
$y(t)$ $={\rm\textbf{C}}h(t)+{\rm\textbf{D}}x(t),$
where $N$ is the state size, ${\rm\textbf{A}}\in\mathbb{R}^{N\times N}$, ${\rm\textbf{B}}\in\mathbb{R}^{N\times 1}$, ${\rm\textbf{C}}\in\mathbb{R}^{1\times N}$, and ${\rm\textbf{D}}\in\mathbb{R}$.
After that, the discretization process is typically adopted to integrate Eq. 1 into practical deep learning algorithms. Specifically, denote $\rm\Delta$ as the timescale parameter to transform the continuous parameters A, B to discrete parameters $\overline{{\rm\textbf{A}}}$, $\overline{{\rm\textbf{B}}}$. The commonly used method for discretization is the zero-order hold (ZOH) rule, which is defined as follows:
$\overline{\rm\textbf{A}}$ $={\rm exp}({\rm{\Delta\textbf{A}}}),$ (2)
$\overline{\rm\textbf{B}}$ $=({\rm{\Delta\textbf{A}}})^{-1}({\rm exp(\textbf{A})}-\textbf{I})\cdot{\rm\Delta\textbf{B}}.$
After the discretization, the discretized version of Eq. 1 with step size $\rm\Delta$ can be rewritten as:
$h_{k}$ $=\overline{\rm\textbf{A}}h_{k-1}+\overline{\rm\textbf{B}}x_{k},$ (3)
$y_{k}$ $={\rm\textbf{C}}h_{k}+{\rm\textbf{D}}x_{k}.$
However, the formulation in Eq. 3 focuses on the LTI system with parameters remaining static for varying inputs. To address this limitation, recent effort [17] attempt to incorporate the selective scan mechanism in which the matrices $\overline{\rm\textbf{B}}$, C and $\rm\Delta$ are derived from the input data. Inspired by the favorable performance, we also employ the selective scan mechanism to facilitate the awareness of the contextual information embedded in the input.
Figure 2: The overall network architecture of our MambaIR, as well as the (a) Residual State-Space Block (RSSB), the (b) Vision State-Space Module (VSSM), and the (c) 2D Selective Scan Module (2D-SSM).
3.2 Overall Architecture
As shown in Fig. 2, our MambaIR consists of three stages: shallow feature extraction, deep feature extraction, and high-quality reconstruction. Given a low-quality (LQ) input image $I_{LQ}\in\mathbb{R}^{H\times W\times 3}$, we first employ a $3\times 3$ convolution layer from the shallow feature extraction to generate the shallow feature $F_{S}\in\mathbb{R}^{H\times W\times C}$, where $H$ and $W$ represent the height and width of the input image, and $C$ is the number of channels. Subsequently, the shallow feature $F_{S}$ undergoes the deep feature extraction stage to acquire the deep feature $F_{D}^{l}\in\mathbb{R}^{H\times W\times C}$ at the $l$-th layer, $l\in\{1,2,\cdots L\}$. This stage is stacked by multiple Residual State-Space Groups (RSSGs), with each RSSG containing several Residue State-Space Blocks (RSSBs). Moreover, an additional convolution layer is introduced at the end of each group to refine features extracted from RSSB. Finally, we use the element-wise sum to obtain the input of the high-quality reconstruction stage $F_{R}=F_{L}+F_{S}$, which is used to reconstruct the high-quality (HQ) output image $I_{HQ}$.
3.3 Residual State-Space Block
The block design in previous Transformer-based image restoration networks [34, 62, 7, 9] mainly follow the Norm $\rightarrow$ Attention $\rightarrow$ Norm $\rightarrow$ MLP flow. Although self-attention and SSM are similar in modeling long-range dependencies, however, we experimentally find that simply replacing Attention with SSM can only obtain sub-optimal results. Therefore, it is promising to tailor a brand-new block structure for Mamba-based restoration networks.
To this end, we propose the Residual State-Space Block (RSSB) to improve the original SSM block with specific restoration priors. Specifically, we consider the local patch recurrence [33] as well as the channel-wise interaction [22] to assist the long-range spatial molding in Mamba. As shown in Fig. 2(a), given the input deep feature $F_{D}^{l}\in\mathbb{R}^{H\times W\times C}$, we first use the Layernorm (LN) followed by the Vision State-Space Module (VSSM) [36] to capture the spatial long-term dependency. Following [8], we also use learnable scale factor $s\in\mathbb{R}^{C}$ in skip connection [8]for better performance:
$Z^{l}=\mathrm{VSSM}(\mathrm{LN}(F_{D}^{l}))+s\cdot F_{D}^{l}.$ (4)
After that, we employ another LayerNorm to normalize the $Z^{l}$ and then use convolution layers to model the spatial local similarity prior. In order to mitigate channel redundancy as well as efficiency, the convolution layer adapts the bottleneck structure, i.e., the channel is first compressed by a factor $\gamma$ to obtain features with the shape $\mathbb{R}^{H\times W\times\frac{C}{\gamma}}$, then we perform channel expansion to recover the original shape. Furthermore, to model the interaction among channels, we introduce the Channel-Attention layer (CA) [22] followed by another tunable scale factor $s^{\prime}\in\mathbb{R}^{C}$ in residual connection to acquire the final output $F_{D}^{l+1}$ of the RSSB. The above process can be formulated as:
$F_{D}^{l+1}=\mathrm{CA(Conv(LN}(Z^{l})))+s^{\prime}\cdot Z^{l}.$ (5)
3.4 Vision State-Space Module
To maintain efficiency, the Transformer-based restoration networks usually divide input into small patches [6] or adopt shifted window attention [34], hindering the interaction at the whole-image level. Motivated by the success of Mamba in long-range modeling with linear complexity, we introduce the Vision State-Space Module to image restoration. Following [36], the input feature $X\in\mathbb{R}^{H\times W\times C}$ will go through two parallel branches. In the first branch, the feature channel is expanded to $\lambda C$ by a linear layer, where $\lambda$ is a pre-defined channel expansion factor, followed by a depth-wise convolution, SiLU [51] activation function, together with the 2D-SSM layer and Layernorm. In the second branch, the features channel is also expanded to $\lambda C$ with a linear layer followed by the SiLU activation function. After that, features from the two branches are aggregated with the Hadamard product. Finally, the channel number is projected back to $C$ to generate output $X_{out}$ with the same shape as input.
$X_{1}=\mathrm{LN(2D\text{-}SSM(SiLU(DWConv(Linear}(X))))),$ (6)
$X_{2}=\mathrm{SiLU(Linear}(X)),$
$X_{out}=\mathrm{Linear}(X_{1}\odot X_{2}),$
where DWConv represents depth-wise convolution, and $\odot$ denotes the Hadamard product.
3.5 2D Selective Scan Module
The standard Mamba [17] causally processes the input data, and thus can only capture information within the scanned part of the data. This property is well suited for NLP tasks that involve a sequential nature but poses significant challenges when transferring to non-causal data such as images. To better utilize the 2D spatial information, we follow [36] and introduce the 2D Selective Scan Module (2D-SSM). As shown in Fig. 2(c), the 2D image feature is flattened into a 1D sequence with scanning along four different directions: top-left to bottom-right, bottom-right to top-left, top-right to bottom-left, and bottom-left to top-right. Then the long-range dependency of each sequence is captured according to the discrete state-space equation in Eq. 3. Finally, all sequences are merged using summation followed by the reshape operation to recover the 2D structure.
3.6 Discussion
Here we would like to discuss the differences between our MambaIR with previous image restoration arts to highlight the contribution of this work. Our MambaIR improves the recent advanced state-space model Mamba with restoration-specific priors, thus allowing the injection of local patch recurrence and channel interaction while maintaining strong long-range capability. Different from Transformer-based methods, the sequential causal property of SSMs exempts us from positional encoding [54]. Meanwhile, the linear complexity with input scale ensures the training and inference efficiency, thus avoiding the shifted window operation which may induce an unsatisfactory effective receptive field.
4 Experiences
4.1 Experimental Settings
Dataset and Evaluation. Following the setup in previous works [34, 62], we conduct experiments on distinct image restoration tasks, including image super-resolution (i.e., classic SR and lightweight SR) and real image denoising. For classic and lightweight image super-resolution, we employ DIV2K [53] and Flickr2K [35] as our training datasets, and use Set5 [3], Set14 [61], B100 [41], Urban100 [23], and Manga109 [42] to evaluate the effectiveness. For real image denoising, we train our model with 320 high-resolution images from SIDD [1] datasets, and use the SIDD test set and DND  [49] dataset for testing. Following [34, 68], we denote the model as MambaIR+ when self-ensemble strategy [35] is used in testing. The performance is evaluated using PSNR and SSIM on the Y channel from the YCbCr color space.
Training Details. In accordance with previous work [34, 7, 62], we perform data augmentation on the training data by applying horizontal flips and random rotations of $90^{\circ},180^{\circ}$, and $270^{\circ}$. Additionally, we crop the original images into $64\times 64$ patches for image SR. For the classic SR, we use the pre-trained weights from the $\times$2 model to initialize those of $\times$3 and $\times$4, and halve the learning rate and total training iterations to reduce training time [35]. To ensure a fair comparison, we adjust the training batch size to 32 for image SR. We employ the Adam [28] as the optimizer for training our MambaIR with $\beta_{1}=0.9,\beta_{2}=0.999$. The initial learning rate is set at $2\times 10^{-4}$ and is halved when the training iteration reaches specific milestones. Our MambaIR model is trained with 8 NVIDIA V100 GPUs.
Table 1: Effects of different design choices of RSSB.
settings Set5 Set14 B100 Urban100 Manga109
(0)baseline 38.57 34.67 32.58 34.15 40.28
(1)+w/o Conv 38.48 34.54 32.56 34.04 40.20
(2)+w/o Conv+CA 38.55 34.64 32.57 34.06 40.14
(3)+replace with MLP 38.55 34.68 32.59 34.22 40.13
4.2 Ablation Study
Effects of different designs of RSSB. As the core component, the RSSB can improve Mamba with restoration-specific priors. In this section, we ablate different components of the RSSB. The PSNR results of ×$\times 2$ classic SR, presented in Tab. 1, indicate that (1) Exploiting the local patch recurrence in the image can assist the long-range spatial modeling of Mamaba, without which will cause a performance drop. (2) Without using local convolution and channel attention, i.e., directly employing off-the-shelf Mamba for restoration, can only obtain sub-optimal results due to the domain gap. (3) Replace Conv+ChanelAttention with MLP, whose structure will be similar to Transformer, also leads to unfavorable results, indicating that although both SSMs and Attention have the global modeling ability, different assisting modules should be considered for further improvements.
4.3 Image Super-Resolution.
Table 2: Quantitative comparison on classic image super-resolution with state-of-the-art methods.he best and the second best results are in red and blue.
Set5 Set14 BSDS100 Urban100 Manga109
Method scale PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
EDSR [35] ×$\times 2$ 38.11 0.9602 33.92 0.9195 32.32 0.9013 32.93 0.9351 39.10 0.9773
RCAN [68] ×$\times 2$ 38.27 0.9614 34.12 0.9216 32.41 0.9027 33.34 0.9384 39.44 0.9786
SAN [11] ×$\times 2$ 38.31 0.9620 34.07 0.9213 32.42 0.9028 33.10 0.9370 39.32 0.9792
HAN [48] ×$\times 2$ 38.27 0.9614 34.16 0.9217 32.41 0.9027 33.35 0.9385 39.46 0.9785
IGNN [70] $\times$2 38.24 0.9613 34.07 0.9217 32.41 0.9025 33.23 0.9383 39.35 0.9786
CSNLN [45] ×$\times 2$ 38.28 0.9616 34.12 0.9223 32.40 0.9024 33.25 0.9386 39.37 0.9785
NLSA [44] ×$\times 2$ 38.34 0.9618 34.08 0.9231 32.43 0.9027 33.42 0.9394 39.59 0.9789
ELAN [66] ×$\times 2$ 38.36 0.9620 34.20 0.9228 32.45 0.9030 33.44 0.9391 39.62 0.9793
IPT [6] ×$\times 2$ 38.37 - 34.43 - 32.48 - 33.76 - - -
SwinIR [34] ×$\times 2$ 38.42 0.9623 34.46 0.9250 32.53 0.9041 33.81 0.9427 39.92 0.9797
SRFormer [71] ×$\times 2$ 38.51 0.9627 34.44 0.9253 32.57 0.9046 34.09 0.9449 40.07 0.9802
MambaIR ×$\times 2$ 38.57 0.9627 34.67 0.9261 32.58 0.9048 34.15 0.9446 40.28 0.9806
MambaIR+ ×$\times 2$ 38.60 0.9628 34.69 0.9260 32.60 0.9048 34.17 0.9443 40.33 0.9806
EDSR [35] $\times$3 34.65 0.9280 30.52 0.8462 29.25 0.8093 28.80 0.8653 34.17 0.9476
RCAN [67] $\times$3 34.74 0.9299 30.65 0.8482 29.32 0.8111 29.09 0.8702 34.44 0.9499
SAN [11] $\times$3 34.75 0.9300 30.59 0.8476 29.33 0.8112 28.93 0.8671 34.30 0.9494
HAN [48] $\times$3 34.75 0.9299 30.67 0.8483 29.32 0.8110 29.10 0.8705 34.48 0.9500
IGNN [70] $\times$3 34.72 0.9298 30.66 0.8484 29.31 0.8105 29.03 0.8696 34.39 0.9496
CSNLN [45] $\times$3 34.74 0.9300 30.66 0.8482 29.33 0.8105 29.13 0.8712 34.45 0.9502
NLSA [44] $\times$3 34.85 0.9306 30.70 0.8485 29.34 0.8117 29.25 0.8726 34.57 0.9508
ELAN [66] $\times$3 34.90 0.9313 30.80 0.8504 29.38 0.8124 29.32 0.8745 34.73 0.9517
IPT [6] $\times$3 34.81 - 30.85 - 29.38 - 29.49 - - -
SwinIR [34] $\times$3 34.97 0.9318 30.93 0.8534 29.46 0.8145 29.75 0.8826 35.12 0.9537
SRformer [71] $\times$3 35.02 0.9323 30.94 0.8540 29.48 0.8156 30.04 0.8865 35.26 0.9543
MambaIR $\times$3 35.08 0.9323 30.99 0.8536 29.51 0.8157 29.93 0.8841 35.43 0.9546
MambaIR+ $\times$3 35.13 0.9326 31.06 0.8541 29.53 0.8162 29.98 0.8838 35.55 0.9549
EDSR [35] ×$\times 4$ 32.46 0.8968 28.80 0.7876 27.71 0.7420 26.64 0.8033 31.02 0.9148
RCAN [68] ×$\times 4$ 32.63 0.9002 28.87 0.7889 27.77 0.7436 26.82 0.8087 31.22 0.9173
SAN [11] ×$\times 4$ 32.64 0.9003 28.92 0.7888 27.78 0.7436 26.79 0.8068 31.18 0.9169
HAN [48] ×$\times 4$ 32.64 0.9002 28.90 0.7890 27.80 0.7442 26.85 0.8094 31.42 0.9177
IGNN [70] $\times$4 32.57 0.8998 28.85 0.7891 27.77 0.7434 26.84 0.8090 31.28 0.9182
CSNLN [45] ×$\times 4$ 32.68 0.9004 28.95 0.7888 27.80 0.7439 27.22 0.8168 31.43 0.9201
NLSA [44] ×$\times 4$ 32.59 0.9000 28.87 0.7891 27.78 0.7444 26.96 0.8109 31.27 0.9184
ELAN [66] ×$\times 4$ 32.75 0.9022 28.96 0.7914 27.83 0.7459 27.13 0.8167 31.68 0.9226
IPT  [6] ×$\times 4$ 32.64 - 29.01 - 27.82 - 27.26 - - -
SwinIR [34] ×$\times 4$ 32.92 0.9044 29.09 0.7950 27.92 0.7489 27.45 0.8254 32.03 0.9260
SRFormer [71] ×$\times 4$ 32.93 0.9041 29.08 0.7953 27.94 0.7502 27.68 0.8311 32.21 0.9271
MambaIR ×$\times 4$ 33.03 0.9046 29.20 0.7961 27.98 0.7503 27.68 0.8287 32.32 0.9272
MambaIR+ ×$\times 4$ 31.13 0.9054 29.25 0.7971 28.01 0.7510 27.80 0.8303 32.48 0.9281
Figure 3: Computational complexity comparison between different methods. We set the standard attention [54] which has a global receptive field as baseline, and denote it as “Full-attn”. We adjust the model to ensure the GPU usage is roughly similar at the beginning, and then scale the input resolution from $48\times 48$ to $84\times 84$.
Figure 4: Qualitative comparison of our MambaIR with CNN- and Transformer-based methods on classic image SR with scale $\times$4.
4.3.1 Classic Image Super-Resolution.
Tab. 2 shows the quantitative results between MambaIR and state-of-the-art super-resolution methods. Thanks to the significant global effective receptive filed, our proposed MambaIR achieves the best performance on almost all five benchmark datasets for all scale factors. For example, our Mamba-based baseline outperforms the Transformer-based benchmark model SwinIR by 0.41dB on Manga109 for ×$\times 2$ scale, demonstrating the prospect of Mamba for image restoration. As for the computational complexity(see Fig. 3), our method is far more efficient than the full-attention baseline [54], and exhibits linear complexity with input resolution which is similar to the efficient attention design such as SwinIR. This observation suggests that out MambaIR has similar scale properties as shifted window attention, and at the same time possesses a global receptive field similar to standard attention. We also give visual comparisons in Fig. 4, and it can be seen that our method can facilitate the reconstruction of sharp edges and natural textures. Moreover, the LAM visualization [20] shown in Fig. 5 demonstrates our MambaIR can utilize more pixels during restoration due to the powerful long-range modeling capability of Mamba.
Figure 5: LAM visualization [20] with CNN and Transformer based methods on $\times$4 SR task. The LAM attribution reflects the importance of each pixel in the input LR image when reconstructing the patch marked with a box. The diffusion index (DI) reflects the range of involved pixels. A higher DI represents a wider range of utilized pixels.
Table 3: Quantitative comparison on lightweight image super-resolution with state-of-the-art methods.
Set5 Set14 BSDS100 Urban100 Manga109
Method scale #param MACs PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
CARN [2] $\times$2 1,592K 222.8G 37.76 0.9590 33.52 0.9166 32.09 0.8978 31.92 0.9256 38.36 0.9765
IMDN [24] $\times$2 694K 158.8G 38.00 0.9605 33.63 0.9177 32.19 0.8996 32.17 0.9283 38.88 0.9774
LAPAR-A [32] $\times$2 548K 171.0G 38.01 0.9605 33.62 0.9183 32.19 0.8999 32.10 0.9283 38.67 0.9772
LatticeNet [39] $\times$2 756K 169.5G 38.15 0.9610 33.78 0.9193 32.25 0.9005 32.43 0.9302 - -
SwinIR-light [34] $\times$2 878K 195.6G 38.14 0.9611 33.86 0.9206 32.31 0.9012 32.76 0.9340 39.12 0.9783
Ours $\times$2 530K 122.1G 38.15 0.9610 33.84 0.9207 32.31 0.9013 32.86 0.9343 39.35 0.9786
CARN [2] $\times$3 1,592K 118.8G 34.29 0.9255 30.29 0.8407 29.06 0.8034 28.06 0.8493 33.50 0.9440
IMDN [24] $\times$3 703K 71.5G 34.36 0.9270 30.32 0.8417 29.09 0.8046 28.17 0.8519 33.61 0.9445
LAPAR-A [32] $\times$3 544K 114.0G 34.36 0.9267 30.34 0.8421 29.11 0.8054 28.15 0.8523 33.51 0.9441
LatticeNet [39] $\times$3 765K 76.3G 34.53 0.9281 30.39 0.8424 29.15 0.8059 28.33 0.8538 - -
SwinIR-light [34] $\times$3 886K 87.2G 34.62 0.9289 30.54 0.8463 29.20 0.8082 28.66 0.8624 33.98 0.9478
Ours $\times$3 538K 55.0G 34.62 0.9286 30.54 0.8459 29.23 0.8083 28.73 0.8635 34.26 0.9482
CARN [2] $\times$4 1,592K 90.9G 32.13 0.8937 28.60 0.7806 27.58 0.7349 26.07 0.7837 30.47 0.9084
IMDN [24] $\times$4 715K 40.9G 32.21 0.8948 28.58 0.7811 27.56 0.7353 26.04 0.7838 30.45 0.9075
LAPAR-A [32] $\times$4 659K 94.0G 32.15 0.8944 28.61 0.7818 27.61 0.7366 26.14 0.7871 30.42 0.9074
LatticeNet [39] $\times$4  [^7] K 43.6G 32.30 0.8962 28.68 0.7830 27.62 0.7367 26.25 0.7873 - -
SwinIR-light [34] $\times$4 897K 49.6G 32.44 0.8976 28.77 0.7858 27.69 0.7406 26.47 0.7980 30.92 0.9151
Ours $\times$4 549K 31.6G 32.41 0.8974 28.76 0.7849 27.70 0.7400 26.53 0.7983 31.05 0.9144
4.3.2 Lightweight Image Super-Resolution.
To demonstrate the scalability of our method, we train the Mamba-light model and compare it with state-of-the-art lightweight image SR methods. Following previous works [71, 39], we also report the number of parameters (#param) and MACs (upscaling a low-resolution image to $1280\times 720$ resolution on all scales). Tab. 3 shows the results. It can be seen that our MambaIR-light outperforms SwinIR-light [34] by up to 0.23dB PSNR on the ×$\times 2$ scale Manga109 dataset with fewer parameters and MACs. The performance results demonstrate the scalability and efficiency of our method.
4.4 Image Denoising
4.4.1 Real Image Denoising.
We further turn to the real image denoising task to evaluate the robustness of our MambaIR when facing real-world degradation. Following [59], we adopt the progressive training strategy for fair comparison. The results are shown in Tab. 4. As one can see, our method achieves comparable performance with existing state-of-the-art models Restormer [57] and outperforms other methods such as Uformer [57] by 0.12dB PSNR on SIDD dataset. The promising results indicate the ability of our method in real image denoising.
Table 4: Quantitative comparison on real image denosing with state-of-the-art methods.
DeamNet [50] MPRNet [60] DAGL [46] Uformer [57] Restormer [59] Ours
Dataset PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
SIDD 39.47 0.957 39.71 0.958 38.94 0.953 39.77 0.959 40.02 0.960 39.89 0.960
DND 39.63 0.953 39.80 0.954 39.77 0.956 39.96 0.956 40.03 0.956 40.04 0.956
5 Limitation and Future Work
While MambaIR appears as a competitive restoration backbone against mainstream Transformer-based methods, it can be further improved with restoration-specific designs. For example, in the proposed method, the local similarity and channel interaction priors are employed to obtain better feature representation. However, other restoration prior such as high-frequency enhancement can be also used. As a pure SSM-based method, our MambaIR does not include the Attention module, therefore how to appropriately combine Attention and Mamba remains an interesting research topic. Moreover, although this work has covered multiple image restoration tasks, some other tasks such as image deblurring and deraining can also be explored in the future. Finally, despite the promising results shown, we would like to point out that the Mamba-based image restoration network is still in its early stages. With the increasing interest in Mamba, it will be promising to study the state-space models for low-level vision.
6 Conclusion
In this work, we explore for the first time the power of the recent advanced state space model i.e., Mamba for image restoration through extensive experiments. By exploiting restoration-specific prior, we propose a brand-new backbone solution, named MambaIR, that can well capture the local patch recurrence and channel interaction while maintaining strong long-range capability and linear complexity, serving as a simple but effective Mamba-based benchmark model for image restoration. Extensive experiments on multiple restoration tasks demonstrate the effectiveness of MambaIR against existing CNN- and Tansformer-based methods.

### v2

@@
 Abstract
-Recent years have witnessed great progress in image restoration thanks to the advancements in modern deep neural networks e.g.
-Convolutional Neural Network and Transformer.
-However, existing restoration backbones are usually limited due to the inherent local reductive bias or quadratic computational complexity.
-Recently, Selective Structured State Space Model e.g., Mamba, has shown great potential for long-range dependencies modeling with linear complexity, but it is still under-explored in low-level computer vision.
-In this work, we introduce a simple but strong benchmark model, named MambaIR, for image restoration.
-In detail, we propose the Residual State Space Block as the core component, which employs convolution and channel attention to enhance the capabilities of the vanilla Mamba.
-In this way, our MambaIR takes advantage of local patch recurrence prior as well as channel interaction to produce restoration-specific feature representation.
-Extensive experiments demonstrate the superiority of our method, for example, MambaIR outperforms Transformer-based baseline SwinIR by up to 0.36dB, using similar computational cost but with a global receptive field.
+Recent years have seen significant advancements in image restoration, largely attributed to the development of modern deep neural networks, such as CNNs and Transformers.
+However, existing restoration backbones often face the dilemma between global receptive fields and efficient computation, hindering their application in practice.
+Recently, the Selective Structured State Space Model, especially the improved version Mamba, has shown great potential for long-range dependency modeling with linear complexity, which offers a way to resolve the above dilemma.
+However, the standard Mamba still faces certain challenges in low-level vision such as local pixel forgetting and channel redundancy.
+In this work, we introduce a simple but effective baseline, named MambaIR, which introduces both local enhancement and channel attention to improve the vanilla Mamba.
+In this way, our MambaIR takes advantage of the local pixel similarity and reduces the channel redundancy.
+Extensive experiments demonstrate the superiority of our method, for example, MambaIR outperforms SwinIR by up to 0.45dB on image SR, using similar computational cost but with a global receptive field.
 Code is available at https://github.com/csguoh/MambaIR.
@@
 Image restoration, aiming to reconstruct a high-quality image from a given low-quality input, is a long-standing problem in computer vision and further has a wide range of sub-problems such as super-resolution, image denoising, etc.
-Since several pioneering works [64, 65, 13], Convolutional Neural Networks (CNNs) have gained great success than traditional model-based method [21, 10], and almost dominate this field in the past few years [30, 56, 35, 67, 11, 63].
+With the introduction of modern deep learning models such as CNNs [70, 14, 37, 76, 12] and Transformers [8, 36, 9, 34, 11], state-of-the-art performance has continued to be refreshed in the past few years.
+To some extent, the increasing performance of deep restoration models largely stems from the increasing network receptive field.
+First, a large receptive field allows the network to capture information from a wider region, enabling it to refer to more pixels to facilitate the reconstruction of the anchor pixel.
+Second, with a larger receptive field, the restoration network can extract higher-level patterns and structures in the image, which can be crucial for some structure preservation tasks such as image denoising.
+Finally, Transformer-based restoration methods which possess larger receptive fields experimentally outperform CNN-based methods, and the recent art [9] also points out that activating more pixels usually leads to better restoration results.
 Figure 1: The Effective Receptive Field (ERF) visualization [40] for EDSR [37], RCAN [76], SwinIR [36], HAT [9], and the proposed MambaIR.
@@
 Only the proposed MambaIR achieves a significant global effective receptive field.
-Despite the promising results, these restoration backbones usually suffer from the inherent problem of CNNs.
-First, the weight-sharing strategy in CNN makes the model remain static for varying inputs, preventing robustness on unseen real-world low-quality images.
-Second, the receptive field of one convolution layer is constrained by the kernel size, which leads to strong local reductive bias.
-Although stacking multiple convolutional layers can theoretically enlarge the receptive field, however as shown in Fig.
-1(a) the effective receptive field [38] of CNN-based restoration methods is still limited.
-Inspired by the success in natural language processing, Transformer [54] has proven superior performance than CNNs in several vision tasks due to weak inductive bias as well as global interaction modeling [4, 14, 37].
-However, taming Transformer for image restoration poses a significant challenge since the self-attention mechanism scales quadratically with the number of tokens.
-Considering that image restoration methods usually set the number of tokens to the image resolution, therefore directly using the attention mechanism will incur a huge computational overhead.
-Although some works make efforts to divide input into small patches [6] or employ shifted window attention [34, 7], these strategies harm the long-term dependency [7] as shown in Fig.
-1(b), and do not essentially address the quadratic complexity problem.
-Recently, structured state-space sequence models (S4), have emerged as an efficient and effective backbone for constructing deep networks [18, 52, 15, 43, 17].
-In particular, the improved S4 with selective mechanism and efficient hardware design, also called Mamba, has been proven to surpass Transformer on tasks requiring long-term dependency modeling such as natural language processing [17].
-More recently, some variants of Mamba have also presented promising results in computer vision [72, 36, 40], such as image classification, biomedical image segmentation, etc.
-However, the power of Mamba in low-level vision is still under-explored, which motivates us to explore the potential of Mamba to enhance the long-range modeling ability for image restoration networks.
-In this work, we propose a simple but effective benchmark restoration model, namely MambaIR, to adapt Mamba for image restoration.
-Following  [34, 7, 62], our MambaIR consists of three stages: shallow feature extraction, deep feature extraction, and high-quality image reconstruction.
-The shallow feature extraction employs a simple convolution layer to extract the shallow feature.
-Then the deep feature extraction is performed with several stacked Residual State Space Blocks (RSSBs).
-As the core component of MambaIR, the RSSB is designed to enhance the capabilities of the original Mamba with image restoration priors.
-Specifically, to tackle the problem that the original Mamba treats each token equally, ignoring the patch recurrence occurring in the neighborhood, we add additional convolution layers to give local pixels more focus.
-We further introduce channel attention layer [22] to supplement channel interaction which is beneficial for image restoration but missed in Mamba.
-In addition, we use the skip connection within each block to offer shortcuts for feature aggregation.
-Finally, the high-quality image reconstruction stage fuses both shallow and deep features to restore the high-quality output image.
-Our main contributions can be summarized as follows:
-  • We are the first to work to adapt state space models for low-level image restoration via extensive experiments to formulate MambaIR, which acts as a simple but effective alternative for CNNs and Transformers.
-  • We propose the Residue State Space Block (RSSB) which can boost the power of the original Mamba with local spatial prior and channel interaction.
-  • Extensive experiments on various tasks demonstrate our MambaIR outperform strong Transformer-based baselines to provide a strong and promising backbone model solution for image restoration.
+Despite possessing many attractive properties, it appears that there exists an inherent choice dilemma between global receptive fields and efficient computation for current image restoration backbones.
+For CNN-based restoration networks [37, 76], although the effective receptive field is limited (as shown in Fig.
+1(a)), it is appropriate for resource-constrained device deployments due to the favorable efficiency of convolution parallel operations.
+By contrast, Transformer-based image restoration methods usually set the number of tokens to the image resolution [36, 8, 9], therefore, despite the global receptive field, directly using the standard Transformer [57] will come at an unacceptable quadratic computational complexity.
+Moreover, employing some efficient attention techniques such as shifted window attention [39] for image restoration, usually comes at the expense of a globally effective receptive field (as shown in  Fig.
+1(b)), and does not intrinsically escape out of the trade-off between a global receptive field and efficient computation.
+Recently, structured state-space sequence models (S4), especially the improved version Mamba, have emerged as an efficient and effective backbone for constructing deep networks [21, 55, 17, 46, 19].
+This development hints at a potential resolution to balancing global receptive field and computational efficiency in image restoration.
+In detail, the discretized state space equations in Mamba can be formalized into a recursive form and can model very long-range dependencies when equipped with specially designed structured reparameterization [20].
+This means that Mamba-based restoration networks can naturally activate more pixels, thus improving the reconstruction quality.
+Furthermore, the parallel scan algorithm [19] renders Mamba to process each token in a parallel fashion, facilitating efficient training on modern hardware such as GPU.
+The above promising properties motivate us to explore the potential of Mamba to achieve efficient long-range modeling for image restoration networks.
+However, the standard Mamba [19], which is designed for 1D sequential data in NLP, is not a natural fit for image restoration scenarios.
+First, since Mamba processes flattened 1D image sequences in a recursive manner, it can result in spatially close pixels being found at very distant locations in the flattened sequences, resulting in the problem of local pixel forgetting.
+Second, due to the requirement to memorize the long sequence dependencies, the number of hidden states in the state space equations is typically large, which can lead to channel redundancy, thus hindering the learning of critical channel representations.
+To address the above challenges, we introduce MambaIR, a simple but very effective benchmark model, to adapt Mamba for image restoration.
+MambaIR is formulated with three principal stages.
+Specifically, the 1)Shallow Feature Extraction stage employs a simple convolution layer to extract the shallow feature.
+Then the 2)Deep Feature Extraction stage performs with several stacked Residual State Space Blocks (RSSBs).
+As the core component of our MambaIR, the RSSB is designed with local convolution to mitigate local pixel forgetting when applying the vanilla Mamba to 2D images, and it is also equipped with channel attention to reduce channel redundancy caused by the excessive hidden state number.
+We also employ the learnable factor to control the skip connection within each RSSB.
+Finally, the 3)High-Quality Image Reconstruction stage aggregates both shallow and deep features to produce a high-quality output image.
+Through possessing both a global effective receptive field as well as linear computational complexity, our MambaIR serves as a new alternative for image restoration backbones.
+In short, our main contributions can be summarized as follows:
+  • We are the first to work to adapt state space models for low-level image restoration via extensive experiments to formulate MambaIR, which acts as a simple but effective alternative for CNN- and Transformer-based methods.
+  • We propose the Residue State Space Block (RSSB) which can boost the power of the standard Mamba with local enhancement and channel redundancy reduction.
+  • Extensive experiments on various tasks demonstrate our MambaIR outperforms other strong baselines to provide a powerful and promising backbone solution for image restoration.
 2 Related Work
 2.1 Image Restoration
-Image restoration has been significantly advanced since the introduction of deep learning by several pioneering works, such as SRCNN [13] for image super-resolution, DnCNN [64] for image denoising and ARCNN [12] for JPEG compression artifact reduction, etc.
+Image restoration has been significantly advanced since the introduction of deep learning by several pioneering works, such as SRCNN [14] for image super-resolution, DnCNN [70] for image denoising, ARCNN [13] for JPEG compression artifact reduction, etc.
 Early attempts usually elaborate CNNs with techniques such as residual connection [29, 67, 6], dense connection [60, 78] and others [31, 62, 18, 76, 12] to improve model representation ability.
-Despite the success, CNN-based restoration methods typically face challenges in effectively modeling global dependencies as well as static weight for varying inputs.
-As Transformers have been proven to be a strong competitor for CNNs in multiple computer vision tasks [4, 14, 37], using Transformer for image restoration appears promising.
-However, it still faces specific challenges from the quadratic computational complexity of the self-attention mechanism [54].
+Despite the success, CNN-based restoration methods typically face challenges in effectively modeling global dependencies.
+As Transformers have been proven to be a strong competitor for CNNs in multiple computer vision tasks [5, 15, 39], using Transformers for image restoration appears promising.
+Despite the global receptive field, it still faces specific challenges from the quadratic computational complexity of the self-attention mechanism [57].
 To address this, IPT [8] divides one image into several small patches and processes each patch independently with self-attention.
-SwinIR [34] further introduces shifted window attention to improve the performance.
-In addition, promising progress continues to be made in designing efficient and effective attention mechanisms for restoration [62, 7, 31, 9, 59, 8].
-Nonetheless, efficient attention design usually comes at the expense of global effective receptive fields, and the issue of quadratic computational complexity is not essentially addressed.
+SwinIR [36] further introduces shifted window attention [39] to improve the performance.
+In addition, progress continues to be made in designing efficient attention mechanisms for restoration [66, 9, 32, 11, 63, 10].
+Nonetheless, efficient attention design usually comes at the expense of global effective receptive fields, and the dilemma of the trade-off between efficient computation and global modeling is not essentially resolved.
 2.2 State Space Models
@@
 More recently, Mamba [19], a data-dependent SSM with selective mechanism and efficient hardware design, outperforms Transformers on natural language and enjoys linear scaling with input length.
-And there are also pioneering works that adopt Mamba to vision tasks such as image classification [36, 72], video understanding [55], biomedical image segmentation [40] and others [25, 47].
-In this work, we explore the potential of Mamba to image restoration with restoration-specific prior to sever as a simple but effective baseline for future work.
+Moreover, there are also pioneering works that adopt Mamba to vision tasks such as image classification [38, 81], video understanding [58], biomedical image segmentation [42] and others [26, 50].
+In this work, we explore the potential of Mamba to image restoration with restoration-specific designs to serve as a simple but effective baseline for future work.
 3 Methodology
-This work aims to introduce the recent advanced state space model (SSM), i.e., Mamba, to low-level image restoration.
-In this section, we begin with a description of the preliminaries of the SSM.
-Then we give an overview of our simple but effective benchmark model MambaIR, followed by a detailed description of the proposed Residule State-Space Block (RSSB).
-Finally, we provide a comprehensive discussion of the difference between our MambaIR and previous methods.
 3.1 Preliminaries
@@
 $\overline{\rm\textbf{B}}$ $=({\rm{\Delta\textbf{A}}})^{-1}({\rm exp(\textbf{A})}-\textbf{I})\cdot{\rm\Delta\textbf{B}}.$
-After the discretization, the discretized version of Eq. 1 with step size $\rm\Delta$ can be rewritten as:
+After the discretization, the discretized version of Eq. 1 with step size $\rm\Delta$ can be rewritten in the following RNN form:
 $h_{k}$ $=\overline{\rm\textbf{A}}h_{k-1}+\overline{\rm\textbf{B}}x_{k},$ (3)
 $y_{k}$ $={\rm\textbf{C}}h_{k}+{\rm\textbf{D}}x_{k}.$
-However, the formulation in Eq.
-3 focuses on the LTI system with parameters remaining static for varying inputs.
-To address this limitation, recent effort [17] attempt to incorporate the selective scan mechanism in which the matrices $\overline{\rm\textbf{B}}$, C and $\rm\Delta$ are derived from the input data.
-Inspired by the favorable performance, we also employ the selective scan mechanism to facilitate the awareness of the contextual information embedded in the input.
-Figure 2: The overall network architecture of our MambaIR, as well as the (a) Residual State-Space Block (RSSB), the (b) Vision State-Space Module (VSSM), and the (c) 2D Selective Scan Module (2D-SSM).
+Furthermore, the Eq. 3 can also be mathematically equivalently transformed into the following CNN form:
+$\overline{\rm\textbf{K}}$ $\triangleq(\mathrm{\textbf{C}}\overline{\mathrm{\textbf{B}}},{\rm{\textbf{C}}}\overline{\rm\textbf{A}}\overline{\rm\textbf{B}},\cdots,{\rm{\textbf{C}}}{\overline{\rm\textbf{A}}}^{L-1}\overline{\rm\textbf{B}}),$ (4)
+y $={\rm\textbf{x}}\circledast\overline{\rm\textbf{K}},$
+[…]
+The recent advanced state-space model, Mamba [19], have further improved $\overline{\rm\textbf{B}}$, C and $\rm\Delta$ to be input-dependent, thus allowing for a dynamic feature representation.
+The intuition of Mamba for image restoration lies in its development on the advantages of S4 model.
+Specifically, Mamba shares the same recursive form of  Eq.
+3, which enables the model to memorize ultra-long sequences so that more pixels can be activated to aid restoration.
+At the same time, the parallel scan algorithm [19] allows Mamba to enjoy the same advantages of parallel processing as Eq.
+4, thus facilitating efficient training.
 3.2 Overall Architecture
@@
 Moreover, an additional convolution layer is introduced at the end of each group to refine features extracted from RSSB.
-Finally, we use the element-wise sum to obtain the input of the high-quality reconstruction stage $F_{R}=F_{L}+F_{S}$, which is used to reconstruct the high-quality (HQ) output image $I_{HQ}$.
+Finally, we use the element-wise sum to obtain the input of the high-quality reconstruction stage $F_{R}=F^{L}_{D}+F_{S}$, which is used to reconstruct the high-quality (HQ) output image $I_{HQ}$.
+Figure 2: The overall network architecture of our MambaIR, as well as the (a) Residual State-Space Block (RSSB), the (b) Vision State-Space Module (VSSM), and the (c) 2D Selective Scan Module (2D-SSM).
 3.3 Residual State-Space Block
-The block design in previous Transformer-based image restoration networks [34, 62, 7, 9] mainly follow the Norm $\rightarrow$ Attention $\rightarrow$ Norm $\rightarrow$ MLP flow.
-Although self-attention and SSM are similar in modeling long-range dependencies, however, we experimentally find that simply replacing Attention with SSM can only obtain sub-optimal results.
+The block design in previous Transformer-based restoration networks [36, 66, 9, 11] mainly follow the Norm $\rightarrow$ Attention $\rightarrow$ Norm $\rightarrow$ MLP flow.
+Although Attention and SSM can both model global dependencies, however, we find these two modules behave differently (see supplementary material for more details) and simply replacing Attention with SSM can only obtain sub-optimal results.
 Therefore, it is promising to tailor a brand-new block structure for Mamba-based restoration networks.
-To this end, we propose the Residual State-Space Block (RSSB) to improve the original SSM block with specific restoration priors.
-Specifically, we consider the local patch recurrence [33] as well as the channel-wise interaction [22] to assist the long-range spatial molding in Mamba.
+To this end, we propose the Residual State-Space Block (RSSB) to adapt the SSM block for restoration.
 As shown in Fig.
 2(a), given the input deep feature $F_{D}^{l}\in\mathbb{R}^{H\times W\times C}$, we first use the LayerNorm (LN) followed by the Vision State-Space Module (VSSM) [38] to capture the spatial long-term dependency.
-Following [8], we also use learnable scale factor $s\in\mathbb{R}^{C}$ in skip connection [8]for better performance:
+Moreover, we also use learnable scale factor $s\in\mathbb{R}^{C}$ to control the information from skip connection:
 $Z^{l}=\mathrm{VSSM}(\mathrm{LN}(F_{D}^{l}))+s\cdot F_{D}^{l}.$ (5)
-After that, we employ another LayerNorm to normalize the $Z^{l}$ and then use convolution layers to model the spatial local similarity prior.
-In order to mitigate channel redundancy as well as efficiency, the convolution layer adapts the bottleneck structure, i.e., the channel is first compressed by a factor $\gamma$ to obtain features with the shape $\mathbb{R}^{H\times W\times\frac{C}{\gamma}}$, then we perform channel expansion to recover the original shape.
-Furthermore, to model the interaction among channels, we introduce the Channel-Attention layer (CA) [22] followed by another tunable scale factor $s^{\prime}\in\mathbb{R}^{C}$ in residual connection to acquire the final output $F_{D}^{l+1}$ of the RSSB.
+Furthermore, since SSMs process flattened feature maps as 1D token sequences, the number of neighborhood pixels in the sequence is greatly influenced by the flattening strategy.
+For example, when employing the four-direction unfolding strategy of  [38], only four nearest neighbors are available to the anchor pixel (see  Fig.
+3(a)), i.e., some spatially close pixels in 2D feature map are instead distant from each other in the 1D token sequence, and this over-distance can lead to local pixel forgetting.
+To this end, we introduce an additional local convolution after VSSM to help restore the neighborhood similarity.
+Specifically, we employ LayerNorm to first normalize the $Z^{l}$ and then use convolution layers to compensate for local features.
+In order to maintain efficiency, the convolution layer adopts the bottleneck structure, i.e., the channel is first compressed by a factor $\gamma$ to obtain features with the shape $\mathbb{R}^{H\times W\times\frac{C}{\gamma}}$, then we perform channel expansion to recover the original shape.
+In addition, SSMs typically introduce a larger number of hidden states to memorize very long-range dependencies, and we visualize the activation results for different channels in  Fig.
+3(b) and find notable channel redundancy.
+To facilitate the expressive power of different channels, we introduce the Channel Attention (CA) [23] to RSSB.
+In this way, SSMs can focus on learning diverse channel representations after which the critical channels are selected by subsequent channel attention, thus avoiding channel redundancy.
+At last, another tunable scale factor $s^{\prime}\in\mathbb{R}^{C}$ is used in residual connection to acquire the final output $F_{D}^{l+1}$ of the RSSB.
 The above process can be formulated as:
 $F_{D}^{l+1}=\mathrm{CA(Conv(LN}(Z^{l})))+s^{\prime}\cdot Z^{l}.$ (6)
+Figure 3: (a) Without using local enhancement will cause spatially close pixels (area in the red box) get forgotten in the flattened 1D sequence due to the long distance.
+(b) We use RELU and global average pooling on the VSSM outputs from the last layer to get the channel activation values.
+Most channels are not activated (i.e., channel redundancy) when channel attention is not used.
 3.4 Vision State-Space Module
@@
 Motivated by the success of Mamba in long-range modeling with linear complexity, we introduce the Vision State-Space Module to image restoration.
+The Vision State-Space Module (VSSM) can capture long-range dependencies with the state space equation, and the architecture of VSSM is shown in  Fig.
+2(b).
 Following [38], the input feature $X\in\mathbb{R}^{H\times W\times C}$ will go through two parallel branches.
@@
 After that, features from the two branches are aggregated with the Hadamard product.
-Finally, the channel number is projected back to $C$ to generate output $X_{out}$ with the same shape as input.
+Finally, the channel number is projected back to $C$ to generate output $X_{out}$ with the same shape as input:
 $X_{1}=\mathrm{LN(2D\text{-}SSM(SiLU(DWConv(Linear}(X))))),$ (7)
@@
 2(c), the 2D image feature is flattened into a 1D sequence with scanning along four different directions: top-left to bottom-right, bottom-right to top-left, top-right to bottom-left, and bottom-left to top-right.
-Then the long-range dependency of each sequence is captured according to the discrete state-space equation in Eq.
-3.
+Then the long-range dependency of each sequence is captured according to the discrete state-space equation.
 Finally, all sequences are merged using summation followed by the reshape operation to recover the 2D structure.
-3.6 Discussion
-Here we would like to discuss the differences between our MambaIR with previous image restoration arts to highlight the contribution of this work.
-Our MambaIR improves the recent advanced state-space model Mamba with restoration-specific priors, thus allowing the injection of local patch recurrence and channel interaction while maintaining strong long-range capability.
-Different from Transformer-based methods, the sequential causal property of SSMs exempts us from positional encoding [54].
-Meanwhile, the linear complexity with input scale ensures the training and inference efficiency, thus avoiding the shifted window operation which may induce an unsatisfactory effective receptive field.
+3.6 Loss Function
+To make a fair comparison with previous works [76, 36, 66], we optimize our MambaIR with $L_{1}$ loss for image SR, which can be formulated as:
+$\mathcal{L}=||I_{HQ}-I_{LQ}||_{1},$ (8)
+where $||\cdot||_{1}$ denotes the $L_{1}$ norm. For image denoising, we utilize the Charbonnier loss [7] with $\epsilon=10^{-3}$:
+$\mathcal{L}=\sqrt{||I_{HQ}-I_{LQ}||^{2}+\epsilon^{2}}.$ (9)
 4 Experiences
@@
 Dataset and Evaluation.
-Following the setup in previous works [34, 62], we conduct experiments on distinct image restoration tasks, including image super-resolution (i.e., classic SR and lightweight SR) and real image denoising.
-For classic and lightweight image super-resolution, we employ DIV2K [53] and Flickr2K [35] as our training datasets, and use Set5 [3], Set14 [61], B100 [41], Urban100 [23], and Manga109 [42] to evaluate the effectiveness.
+Following the setup in previous works [36, 66], we conduct experiments on various image restoration tasks, including image super-resolution (i.e., classic SR, lightweight SR, and real SR) and image denoising (i.e., Gaussian color image denoising and real-world denoising).
+We employ DIV2K [56] and Flickr2K [37] to train classic SR models and use only DIV2K to train lightweight SR models.
+Moreover, we use Set5 [4], Set14 [65], B100 [44], Urban100 [24], and Manga109 [45] to evaluate the effectiveness of different SR methods.
+For gaussian color image denoising, we utilize DIV2K [56], Flickr2K [37], BSD500 [3], and WED [43] as our training datasets.
+Our testing datasets for guassian color image denoising includes BSD68 [44], Kodak24 [16], McMaster [74], and Urban100 [24].
 For real image denoising, we train our model with 320 high-resolution images from SIDD [1] datasets, and use the SIDD test set and DND  [52] dataset for testing.
@@
 The performance is evaluated using PSNR and SSIM on the Y channel from the YCbCr color space.
+In addition, we also include other restoration tasks such as JPEG compression artifact reduction, more experimental results can be seen in the supplementary material.
+Table 1: Ablation experiments for different design choices of RSSB.
+settings #param MACs Set5 Set14 B100 Urban100 Manga109
+(0)baseline 16.7M 439G 38.57 34.67 32.58 34.15 40.28
+(1)+w/o Conv 11.9M 316G 38.48 34.54 32.56 34.04 40.20
+(2)+w/o Conv+CA 11.8M 315G 38.55 34.64 32.57 34.06 40.14
+(3)+replace with MLP 14.3M 379G 38.55 34.68 32.59 34.22 40.13
 Training Details.
-In accordance with previous work [34, 7, 62], we perform data augmentation on the training data by applying horizontal flips and random rotations of $90^{\circ},180^{\circ}$, and $270^{\circ}$.
-Additionally, we crop the original images into $64\times 64$ patches for image SR.
-For the classic SR, we use the pre-trained weights from the $\times$2 model to initialize those of $\times$3 and $\times$4, and halve the learning rate and total training iterations to reduce training time [35].
-To ensure a fair comparison, we adjust the training batch size to 32 for image SR.
+In accordance with previous works [36, 9, 66], we perform data augmentation by applying horizontal flips and random rotations of $90^{\circ},180^{\circ}$, and $270^{\circ}$.
+Additionally, we crop the original images into $64\times 64$ patches for image SR and $128\times 128$ patches for image denoising during training.
+For image SR, we use the pre-trained weights from the $\times$2 model to initialize those of $\times$3 and $\times$4 and halve the learning rate and total training iterations to reduce training time [37].
+To ensure a fair comparison, we adjust the training batch size to 32 for image SR and 16 for image denoising.
 We employ the Adam [30] as the optimizer for training our MambaIR with $\beta_{1}=0.9,\beta_{2}=0.999$.
@@
 Our MambaIR model is trained with 8 NVIDIA V100 GPUs.
-Table 1: Effects of different design choices of RSSB.
-settings Set5 Set14 B100 Urban100 Manga109
-(0)baseline 38.57 34.67 32.58 34.15 40.28
-(1)+w/o Conv 38.48 34.54 32.56 34.04 40.20
-(2)+w/o Conv+CA 38.55 34.64 32.57 34.06 40.14
-(3)+replace with MLP 38.55 34.68 32.59 34.22 40.13
 4.2 Ablation Study
@@
 In this section, we ablate different components of the RSSB.
-The PSNR results of ×$\times 2$ classic SR, presented in Tab.
-1, indicate that (1) Exploiting the local patch recurrence in the image can assist the long-range spatial modeling of Mamaba, without which will cause a performance drop.
-(2) Without using local convolution and channel attention, i.e., directly employing off-the-shelf Mamba for restoration, can only obtain sub-optimal results due to the domain gap.
-(3) Replace Conv+ChanelAttention with MLP, whose structure will be similar to Transformer, also leads to unfavorable results, indicating that although both SSMs and Attention have the global modeling ability, different assisting modules should be considered for further improvements.
-4.3 Image Super-Resolution.
-Table 2: Quantitative comparison on classic image super-resolution with state-of-the-art methods.he best and the second best results are in red and blue.
+The results, presented in Tab.
+1, indicate that (1) directly using the standard Mamba to process flattened sequences can lead to local pixel forgetting, and the utilization of simple convolution layers can effectively enhance the local interaction.
+(2) Without using additional convolution and channel attention, i.e., directly employing off-the-shelf Mamba for restoration, can only obtain sub-optimal results, which also supports our previous analysis.
+(3) Replacing Conv+ChanelAttention with MLP, whose resulted structure will be similar to Transformer, also leads to unfavorable results, indicating that although both SSMs and Attention have the global modeling ability, the behavior of these two modules is different and thus accustomed block structure should be considered for further improvements.
+Figure 4: Qualitative comparison of our MambaIR with CNN and Transformer based methods on classic image SR with scale $\times$4.
+Table 2: Ablation experiments for different scan modes in VSSM.
+scan mode #param MACs Set5 Set14 B100 Urban100 Manga109
+one-direction 16.7M 439G 38.53 34.63 32.58 34.06 40.31
+two-direction 16.7M 439G 38.56 34.60 32.56 33.96 40.14
+baseline 16.7M 439G 38.57 34.67 32.58 34.15 40.28
+Table 3: Quantitative comparison on classic image super-resolution with state-of-the-art methods. The best and the second best results are in red and blue.
 Set5 Set14 BSDS100 Urban100 Manga109
@@
 MambaIR+ ×$\times 4$ 31.13 0.9054 29.25 0.7971 28.01 0.7510 27.80 0.8303 32.48 0.9281
-Figure 3: Computational complexity comparison between different methods.
-We set the standard attention [54] which has a global receptive field as baseline, and denote it as “Full-attn”.
-We adjust the model to ensure the GPU usage is roughly similar at the beginning, and then scale the input resolution from $48\times 48$ to $84\times 84$.
-Figure 4: Qualitative comparison of our MambaIR with CNN- and Transformer-based methods on classic image SR with scale $\times$4.
-4.3.1 Classic Image Super-Resolution.
+Effects of Different Scan Modes in VSSM.
+To allow Mamba to process 2D images, the feature map needs to be flattened before being iterated by the state-space equation.
+Therefore, the unfolding strategy is particularly important.
+In this work, we follow  [38] which uses scans in four different directions to generate scanned sequences.
+Here, we ablate different scan modes to study the effects, the results are shown in Tab.
+2.
+Compared with one-direction(top-left to bottom-right) and two-direction(top-left to bottom-right, bottom-right to top-left), using four directions of scanning allows the anchor pixel to perceive a wider range of neighborhoods, thus achieving better results.
+Surprisingly, using more scanning directions brings an almost negligible increase in the number of parameters and computational complexity.
+This benefit may facilitate future exploration of more effective scanning mechanisms for the Mamba-based image restoration networks.
+We also include other ablation experiments, such as the layer number of RSSBs, please see supplementary material for more analysis.
+4.3 Comparison on Image Super-Resolution
+Classic Image Super-Resolution.
 Tab.
 3 shows the quantitative results between MambaIR and state-of-the-art super-resolution methods.
-Thanks to the significant global effective receptive filed, our proposed MambaIR achieves the best performance on almost all five benchmark datasets for all scale factors.
+Thanks to the significant global receptive field, our proposed MambaIR achieves the best performance on almost all five benchmark datasets for all scale factors.
 For example, our Mamba-based baseline outperforms the Transformer-based benchmark model SwinIR by 0.41dB on Manga109 for ×$\times 2$ scale, demonstrating the prospect of Mamba for image restoration.
-As for the computational complexity(see Fig.
-3), our method is far more efficient than the full-attention baseline [54], and exhibits linear complexity with input resolution which is similar to the efficient attention design such as SwinIR.
-This observation suggests that out MambaIR has similar scale properties as shifted window attention, and at the same time possesses a global receptive field similar to standard attention.
 We also give visual comparisons in Fig.
 4, and it can be seen that our method can facilitate the reconstruction of sharp edges and natural textures.
-Moreover, the LAM visualization [20] shown in Fig.
-5 demonstrates our MambaIR can utilize more pixels during restoration due to the powerful long-range modeling capability of Mamba.
-Figure 5: LAM visualization [20] with CNN and Transformer based methods on $\times$4 SR task.
-The LAM attribution reflects the importance of each pixel in the input LR image when reconstructing the patch marked with a box.
-The diffusion index (DI) reflects the range of involved pixels.
-A higher DI represents a wider range of utilized pixels.
+Table 4: Model size comparisons. We compare the complexity on the ×$\times 4$ classic SR model of different methods, the output size is set to $3\times 640\times 640$ for MACs.
+Method EDSR [37] RCAN  [76] SRFBN  [35] HAN  [51] CSNLN  [48] SwinIR  [36] MambaIR (Ours)
+Params (M) 43.1 15.6 3.6 16.1 7.2 11.9 16.7
+Mult-Adds $(\mathrm{G})$ 1,286 407 498 420 103,640 336 439
+PSNR on Urban100 (dB) 26.64 26.82 26.60 26.85 27.22 27.45 27.68
+PSNR on Manga109 (dB) 31.02 31.22 31.15 31.42 31.43 32.03 32.32
+Figure 5: Computational complexity comparison with different input scales.
+We set the standard attention [57] which has a global receptive field as baseline, and denote it as "Full-attn".
+We adjust the model to ensure the GPU usage is roughly similar at the beginning, and then scale the input resolution from $48\times 48$ to $84\times 84$.
+Model Complexity Comparison.
+Tab.
+4 provides the comparison of different methods on parameters and MACs.
+Our MambaIR has comparable parameters and MACs compared with SwinIR, but can achieve a global receptive field and better PSNR performance.
+We further give comparison results of computational complexity with input sizes in  Fig.
+5.
+As one can see, our method is far more efficient than the full-attention baseline [57] and exhibits linear complexity with input resolution which is similar to the efficient attention techniques such as SwinIR.
+These observations above suggest that out MambaIR has similar scale properties as shifted window attention, while possessing a global receptive field similar to standard full attention.
+Lightweight Image Super-Resolution.
+To demonstrate the scalability of our method, we train the Mamba-light model and compare it with state-of-the-art lightweight image SR methods.
+Following previous works [80, 41], we also report the number of parameters (#param) and MACs (upscaling a low-resolution image to $1280\times 720$ resolution).
+Tab.
+5 shows the results.
+It can be seen that our MambaIR-light outperforms SwinIR-light [36] by up to 0.17dB PSNR on the ×$\times 4$ scale Manga109 dataset with similar parameters and MACs.
+The performance results demonstrate the scalability and efficiency of our method.
 Table 5: Quantitative comparison on lightweight image super-resolution with state-of-the-art methods.
@@
 SwinIR-light [36] $\times$2 878K 195.6G 38.14 0.9611 33.86 0.9206 32.31 0.9012 32.76 0.9340 39.12 0.9783
-Ours $\times$2 530K 122.1G 38.15 0.9610 33.84 0.9207 32.31 0.9013 32.86 0.9343 39.35 0.9786
+SRFormer-light [80] $\times$2 853K 236G 38.23 0.9613 33.94 0.9209 32.36 0.9019 32.91 0.9353 39.28 0.9785
+Ours $\times$2 859K 198.1G 38.16 0.9610 34.00 0.9212 32.34 0.9017 32.92 0.9356 39.31 0.9779
 CARN [2] $\times$3 1,592K 118.8G 34.29 0.9255 30.29 0.8407 29.06 0.8034 28.06 0.8493 33.50 0.9440
@@
 SwinIR-light [36] $\times$3 886K 87.2G 34.62 0.9289 30.54 0.8463 29.20 0.8082 28.66 0.8624 33.98 0.9478
-Ours $\times$3 538K 55.0G 34.62 0.9286 30.54 0.8459 29.23 0.8083 28.73 0.8635 34.26 0.9482
+SRFormer-light [80] $\times$3 861K 105G 34.67 0.9296 30.57 0.8469 29.26 0.8099 28.81 0.8655 34.19 0.9489
+Ours $\times$3 867K 88.7G 34.72 0.9296 30.63 0.8475 29.29 0.8099 29.00 0.8689 34.39 0.9495
 CARN [2] $\times$4 1,592K 90.9G 32.13 0.8937 28.60 0.7806 27.58 0.7349 26.07 0.7837 30.47 0.9084
@@
 SwinIR-light [36] $\times$4 897K 49.6G 32.44 0.8976 28.77 0.7858 27.69 0.7406 26.47 0.7980 30.92 0.9151
-Ours $\times$4 549K 31.6G 32.41 0.8974 28.76 0.7849 27.70 0.7400 26.53 0.7983 31.05 0.9144
-4.3.2 Lightweight Image Super-Resolution.
-To demonstrate the scalability of our method, we train the Mamba-light model and compare it with state-of-the-art lightweight image SR methods.
-Following previous works [71, 39], we also report the number of parameters (#param) and MACs (upscaling a low-resolution image to $1280\times 720$ resolution on all scales).
-Tab.
-3 shows the results.
-It can be seen that our MambaIR-light outperforms SwinIR-light [34] by up to 0.23dB PSNR on the ×$\times 2$ scale Manga109 dataset with fewer parameters and MACs.
-The performance results demonstrate the scalability and efficiency of our method.
-4.4 Image Denoising
-4.4.1 Real Image Denoising.
-We further turn to the real image denoising task to evaluate the robustness of our MambaIR when facing real-world degradation.
+SRFormer-light [80] $\times$4 873K 62.8G 32.51 0.8988 28.82 0.7872 27.73 0.7422 26.67 0.8032 31.17 0.9165
+Ours $\times$4 879K 50.6G 32.51 0.8993 28.85 0.7876 27.75 0.7423 26.75 0.8051 31.26 0.9175
+Real-world Image Super-resolution.
+We also investigate the performance of the network for real-world image restoration.
+We follow the training protocol in  [9] to train our MambaIR-real model.
+Since there are no ground-truth images for this task, only the visual comparison is given in Fig.
+6.
+Compared with the other methods, our MambaIR exhibits a notable advancement in resolving fine details and texture preservation, demonstrating the robustness of our method.
+Figure 6: Qualitative comparison with RealSR [27], ESRGAN [60], BSRGAN [69], Real-ESRGAN [59], and SwinIR [36] on real image super-resolution with scale $\times$4.
+4.4 Comparison on Image Denoising
+Gaussian Color Image Denoising.
+The results of gaussian color image denoising are shown in  Tab.
+6.
+Following [67, 70], the compared noise levels include 15, 25 and 50.
+As one can see, our model achieves better performance than all compared methods.
+In particular, it surpasses the SwinIR [36] by up to 0.25dB with $\sigma$=50 on the Urban100 dataset.
+We also give a visual comparison in Fig.
+7.
+Thanks to the global receptive field, our MambaIR can achieve better structure preservation, leading to clearer edges and natural shapes.
+Table 6: Quantitative comparison on gaussian color image denoising with state-of-the-art methods.
+Method BSD68 Kodak24 McMaster Urban100
+$\sigma$=15 $\sigma$=25 $\sigma$=50 $\sigma$=15 $\sigma$=25 $\sigma$=50 $\sigma$=15 $\sigma$=25 $\sigma$=50 $\sigma$=15 $\sigma$=25 $\sigma$=50
+IRCNN [72] 33.86 31.16 27.86 34.69 32.18 28.93 34.58 32.18 28.91 33.78 31.20 27.70
+FFDNet [73] 33.87 31.21 27.96 34.63 32.13 28.98 34.66 32.35 29.18 33.83 31.40 28.05
+DnCNN [71] 33.90 31.24 27.95 34.60 32.14 28.95 33.45 31.52 28.62 32.98 30.81 27.59
+DRUNet [68] 34.30 31.69 28.51 35.31 32.89 29.86 35.40 33.14 30.08 34.81 32.60 29.61
+SwinIR [36] 34.42 31.78 28.56 35.34 32.89 29.79 35.61 33.20 30.22 35.13 32.90 29.82
+Restormer [63] 34.40 31.79 28.60 35.47 33.04 30.01 35.61 33.34 30.30 35.13 32.96 30.02
+MambaIR 34.43 31.80 28.61 35.34 32.91 29.85 35.62 33.35 30.31 35.17 32.99 30.07
+MambaIR+ 34.44 31.82 28.64 35.35 32.92 29.87 35.63 33.36 30.32 35.17 32.99 30.06
+Figure 7: Qualitative comparison of our MambaIR with other methods on color image denoising task with noise level level $\sigma$=50.
+Real Image Denoising.
+We further turn to the real image denoising task to evaluate the robustness of MambaIR when facing real-world degradation.
 Following [63], we adopt the progressive training strategy for fair comparison.
-The results are shown in Tab.
-4.
-As one can see, our method achieves comparable performance with existing state-of-the-art models Restormer [57] and outperforms other methods such as Uformer [57] by 0.12dB PSNR on SIDD dataset.
+The results, shown in Tab.
+7, suggest that our method achieves comparable performance with existing state-of-the-art models Restormer [61] and outperforms other methods such as Uformer [61] by 0.12dB PSNR on SIDD dataset.
 The promising results indicate the ability of our method in real image denoising.
@@
 DND 39.63 0.953 39.80 0.954 39.77 0.956 39.96 0.956 40.03 0.956 40.04 0.956
-5 Limitation and Future Work
-While MambaIR appears as a competitive restoration backbone against mainstream Transformer-based methods, it can be further improved with restoration-specific designs.
-For example, in the proposed method, the local similarity and channel interaction priors are employed to obtain better feature representation.
-However, other restoration prior such as high-frequency enhancement can be also used.
-As a pure SSM-based method, our MambaIR does not include the Attention module, therefore how to appropriately combine Attention and Mamba remains an interesting research topic.
-Moreover, although this work has covered multiple image restoration tasks, some other tasks such as image deblurring and deraining can also be explored in the future.
-Finally, despite the promising results shown, we would like to point out that the Mamba-based image restoration network is still in its early stages.
-With the increasing interest in Mamba, it will be promising to study the state-space models for low-level vision.
 5 Conclusion
-In this work, we explore for the first time the power of the recent advanced state space model i.e., Mamba for image restoration through extensive experiments.
-By exploiting restoration-specific prior, we propose a brand-new backbone solution, named MambaIR, that can well capture the local patch recurrence and channel interaction while maintaining strong long-range capability and linear complexity, serving as a simple but effective Mamba-based benchmark model for image restoration.
-Extensive experiments on multiple restoration tasks demonstrate the effectiveness of MambaIR against existing CNN- and Tansformer-based methods.
+In this work, we explore for the first time the power of the recent advanced state space model, i.e., Mamba, for image restoration, to help resolve the dilemma of trade-off between efficient computation and global effective receptive field.
+Specifically, we introduce the local enhancement to mitigate the neighborhood pixel forgetting problem from the flattening strategy and propose channel attention to reduce channel redundancy.
+Extensive experiments on multiple restoration tasks demonstrate our MambaIR serves as a simple but effective state-space model for image restoration.

### v3

@@
-
+3ByteDance Inc.   4Aitist.ai   5Peng Cheng Laboratory
 Abstract
@@
 Second, with a larger receptive field, the restoration network can extract higher-level patterns and structures in the image, which can be crucial for some structure preservation tasks such as image denoising.
-Finally, Transformer-based restoration methods which possess larger receptive fields experimentally outperform CNN-based methods, and the recent art [9] also points out that activating more pixels usually leads to better restoration results.
-Figure 1: The Effective Receptive Field (ERF) visualization [40] for EDSR [37], RCAN [76], SwinIR [36], HAT [9], and the proposed MambaIR.
+Finally, Transformer-based restoration methods which possess larger receptive fields experimentally outperform CNN-based methods, and the recent work [10] also points out that activating more pixels usually leads to better restoration results.
+Figure 1: The Effective Receptive Field (ERF) visualization [46, 14] for EDSR [42], RCAN [88], SwinIR [41], HAT [10], and the proposed MambaIR.
 A larger ERF is indicated by a more extensively distributed dark area.
-Only the proposed MambaIR achieves a significant global effective receptive field.
+The proposed MambaIR achieves a significant global effective receptive field.
 Despite possessing many attractive properties, it appears that there exists an inherent choice dilemma between global receptive fields and efficient computation for current image restoration backbones.
@@
 1(a)), it is appropriate for resource-constrained device deployments due to the favorable efficiency of convolution parallel operations.
-By contrast, Transformer-based image restoration methods usually set the number of tokens to the image resolution [36, 8, 9], therefore, despite the global receptive field, directly using the standard Transformer [57] will come at an unacceptable quadratic computational complexity.
+By contrast, the Transformer-based image restoration methods usually set the number of tokens to the image resolution [41, 8, 10], therefore, despite the global receptive field, directly using the standard Transformer [65] will come at an unacceptable quadratic computational complexity.
 Moreover, employing some efficient attention techniques such as shifted window attention [45] for image restoration, usually comes at the expense of a globally effective receptive field (as shown in  Fig.
@@
 Recently, structured state-space sequence models (S4), especially the improved version Mamba, have emerged as an efficient and effective backbone for constructing deep networks [24, 62, 18, 52, 22].
-This development hints at a potential resolution to balancing global receptive field and computational efficiency in image restoration.
+This development hints at a potential solution to balancing global receptive field and computational efficiency in image restoration.
 In detail, the discretized state space equations in Mamba can be formalized into a recursive form and can model very long-range dependencies when equipped with specially designed structured reparameterization [23].
@@
 In short, our main contributions can be summarized as follows:
-  • We are the first to work to adapt state space models for low-level image restoration via extensive experiments to formulate MambaIR, which acts as a simple but effective alternative for CNN- and Transformer-based methods.
+  • We are the first work to adapt state space models for low-level image restoration via extensive experiments to formulate MambaIR, which acts as a simple but effective alternative for CNN- and Transformer-based methods.
   • We propose the Residue State Space Block (RSSB) which can boost the power of the standard Mamba with local enhancement and channel redundancy reduction.
@@
 Image restoration has been significantly advanced since the introduction of deep learning by several pioneering works, such as SRCNN [16] for image super-resolution, DnCNN [81] for image denoising, ARCNN [15] for JPEG compression artifact reduction, etc.
-Early attempts usually elaborate CNNs with techniques such as residual connection [29, 67, 6], dense connection [60, 78] and others [31, 62, 18, 76, 12] to improve model representation ability.
+Early attempts usually elaborate CNNs with techniques such as residual connection [34, 6], dense connection [68, 89] and others [36, 70, 19, 13] to improve model representation ability.
 Despite the success, CNN-based restoration methods typically face challenges in effectively modeling global dependencies.
-As Transformers have been proven to be a strong competitor for CNNs in multiple computer vision tasks [5, 15, 39], using Transformers for image restoration appears promising.
-Despite the global receptive field, it still faces specific challenges from the quadratic computational complexity of the self-attention mechanism [57].
+As transformer have proven its effectiveness in multiple tasks, such as time series [43], 3D cloud [77, 75], and multi-modal [20, 4, 21, 86], using transformer for image restoration appears promising.
+Despite the global receptive field, transformer still faces specific challenges from the quadratic computational complexity of the self-attention [65].
 To address this, IPT [8] divides one image into several small patches and processes each patch independently with self-attention.
 SwinIR [41] further introduces shifted window attention [45] to improve the performance.
-In addition, progress continues to be made in designing efficient attention mechanisms for restoration [66, 9, 32, 11, 63, 10].
-Nonetheless, efficient attention design usually comes at the expense of global effective receptive fields, and the dilemma of the trade-off between efficient computation and global modeling is not essentially resolved.
+In addition, progress continues to be made in designing efficient attention for restoration [78, 10, 38, 12, 72, 11, 63, 26, 85, 9].
+Nonetheless, efficient attention design usually comes at the expense of global receptive fields, and the dilemma of the trade-off between efficient computation and global modeling is not essentially resolved.
 2.2 State Space Models
@@
 More recently, Mamba [22], a data-dependent SSM with selective mechanism and efficient hardware design, outperforms Transformers on natural language and enjoys linear scaling with input length.
-Moreover, there are also pioneering works that adopt Mamba to vision tasks such as image classification [38, 81], video understanding [58], biomedical image segmentation [42] and others [26, 50].
+Moreover, there are also pioneering works that adopt Mamba to vision tasks such as image classification [44, 92], video understanding [66, 37], biomedical image segmentation [48, 71] and others [31, 56, 28, 76, 59].
 In this work, we explore the potential of Mamba to image restoration with restoration-specific designs to serve as a simple but effective baseline for future work.
@@
 Dataset and Evaluation.
-Following the setup in previous works [36, 66], we conduct experiments on various image restoration tasks, including image super-resolution (i.e., classic SR, lightweight SR, and real SR) and image denoising (i.e., Gaussian color image denoising and real-world denoising).
-We employ DIV2K [56] and Flickr2K [37] to train classic SR models and use only DIV2K to train lightweight SR models.
+Following the setup in previous works [41, 78], we conduct experiments on various image restoration tasks, including image super-resolution (i.e., classic SR, lightweight SR, real SR) and image denoising (i.e., Gaussian color image denoising and real-world denoising), and JPEG compression artifact reduction (JPEG CAR).
+We employ DIV2K [64] and Flickr2K [42] to train classic SR models and use DIV2K only to train lightweight SR models.
 Moreover, we use Set5 [5], Set14 [74], B100 [50], Urban100 [29], and Manga109 [51] to evaluate the effectiveness of different SR methods.
@@
 The performance is evaluated using PSNR and SSIM on the Y channel from the YCbCr color space.
-In addition, we also include other restoration tasks such as JPEG compression artifact reduction, more experimental results can be seen in the supplementary material.
-Table 1: Ablation experiments for different design choices of RSSB.
-settings #param MACs Set5 Set14 B100 Urban100 Manga109
-(0)baseline 16.7M 439G 38.57 34.67 32.58 34.15 40.28
-(1)+w/o Conv 11.9M 316G 38.48 34.54 32.56 34.04 40.20
-(2)+w/o Conv+CA 11.8M 315G 38.55 34.64 32.57 34.06 40.14
-(3)+replace with MLP 14.3M 379G 38.55 34.68 32.59 34.22 40.13
+Due to page limit, the results of JPEG CAR are shown in the supplementary material.
 Training Details.
@@
 The results, presented in Tab.
-1, indicate that (1) directly using the standard Mamba to process flattened sequences can lead to local pixel forgetting, and the utilization of simple convolution layers can effectively enhance the local interaction.
+1, indicate that (1) applying 1D scan on flattened images can lead to local pixel forgetting, and the utilization of simple convolution layers can effectively enhance the local interaction.
 (2) Without using additional convolution and channel attention, i.e., directly employing off-the-shelf Mamba for restoration, can only obtain sub-optimal results, which also supports our previous analysis.
 (3) Replacing Conv+ChanelAttention with MLP, whose resulted structure will be similar to Transformer, also leads to unfavorable results, indicating that although both SSMs and Attention have the global modeling ability, the behavior of these two modules is different and thus accustomed block structure should be considered for further improvements.
+Table 1: Ablation experiments for different design choices of RSSB.
+settings Set5 Set14 Urban100
+(1)remove Conv 38.48 34.54 34.04
+(2)remove Conv+CA 38.55 34.64 34.06
+(3)replace with MLP 38.55 34.68 34.22
+Table 2: Ablation experiments for different scan modes in VSSM.
+scan mode Set5 Set14 Urban100
+one-direction 38.53 34.63 34.06
+two-direction 38.56 34.60 33.96
+baseline 38.57 34.67 34.15
 Figure 4: Qualitative comparison of our MambaIR with CNN and Transformer based methods on classic image SR with scale $\times$4.
-Table 2: Ablation experiments for different scan modes in VSSM.
-scan mode #param MACs Set5 Set14 B100 Urban100 Manga109
-one-direction 16.7M 439G 38.53 34.63 32.58 34.06 40.31
-two-direction 16.7M 439G 38.56 34.60 32.56 33.96 40.14
-baseline 16.7M 439G 38.57 34.67 32.58 34.15 40.28
 Table 3: Quantitative comparison on classic image super-resolution with state-of-the-art methods. The best and the second best results are in red and blue.
@@
 MambaIR ×$\times 4$ 33.03 0.9046 29.20 0.7961 27.98 0.7503 27.68 0.8287 32.32 0.9272
-MambaIR+ ×$\times 4$ 31.13 0.9054 29.25 0.7971 28.01 0.7510 27.80 0.8303 32.48 0.9281
+MambaIR+ ×$\times 4$ 33.13 0.9054 29.25 0.7971 28.01 0.7510 27.80 0.8303 32.48 0.9281
 Effects of Different Scan Modes in VSSM.
@@
 Compared with one-direction (top-left to bottom-right) and two-direction (top-left to bottom-right, bottom-right to top-left), using four directions of scanning allows the anchor pixel to perceive a wider range of neighborhoods, thus achieving better results.
-Surprisingly, using more scanning directions brings an almost negligible increase in the number of parameters and computational complexity.
-This benefit may facilitate future exploration of more effective scanning mechanisms for the Mamba-based image restoration networks.
 We also include other ablation experiments, such as the layer number of RSSBs, please see supplementary material for more analysis.
@@
 We also give visual comparisons in Fig.
-4, and it can be seen that our method can facilitate the reconstruction of sharp edges and natural textures.
-Table 4: Model size comparisons. We compare the complexity on the ×$\times 4$ classic SR model of different methods, the output size is set to $3\times 640\times 640$ for MACs.
-Method EDSR [37] RCAN  [76] SRFBN  [35] HAN  [51] CSNLN  [48] SwinIR  [36] MambaIR (Ours)
-Params (M) 43.1 15.6 3.6 16.1 7.2 11.9 16.7
-Mult-Adds $(\mathrm{G})$ 1,286 407 498 420 103,640 336 439
-PSNR on Urban100 (dB) 26.64 26.82 26.60 26.85 27.22 27.45 27.68
-PSNR on Manga109 (dB) 31.02 31.22 31.15 31.42 31.43 32.03 32.32
+4, and our method can facilitate the reconstruction of sharp edges and natural textures.
 Figure 5: Computational complexity comparison with different input scales.
@@
 Model Complexity Comparison.
-Tab.
-4 provides the comparison of different methods on parameters and MACs.
-Our MambaIR has comparable parameters and MACs compared with SwinIR, but can achieve a global receptive field and better PSNR performance.
-We further give comparison results of computational complexity with input sizes in  Fig.
+We give comparison results of computational complexity with varying input sizes in  Fig.
 5.
@@
 Lightweight Image Super-Resolution.
-To demonstrate the scalability of our method, we train the Mamba-light model and compare it with state-of-the-art lightweight image SR methods.
+To demonstrate the scalability of our method, we train the MambaIR-light model and compare it with state-of-the-art lightweight image SR methods.
 Following previous works [91, 47], we also report the number of parameters (#param) and MACs (upscaling a low-resolution image to $1280\times 720$ resolution).
@@
 4 shows the results.
-It can be seen that our MambaIR-light outperforms SwinIR-light [36] by up to 0.17dB PSNR on the ×$\times 4$ scale Manga109 dataset with similar parameters and MACs.
+It can be seen that our MambaIR-light outperforms SwinIR-light [41] by up to 0.34dB PSNR on the ×$\times 4$ scale Manga109 dataset with similar parameters and MACs.
 The performance results demonstrate the scalability and efficiency of our method.
@@
 LAPAR-A [39] $\times$2 548K 171.0G 38.01 0.9605 33.62 0.9183 32.19 0.8999 32.10 0.9283 38.67 0.9772
-LatticeNet [41] $\times$2 756K 169.5G 38.15 0.9610 33.78 0.9193 32.25 0.9005 32.43 0.9302 - -
-SwinIR-light [36] $\times$2 878K 195.6G 38.14 0.9611 33.86 0.9206 32.31 0.9012 32.76 0.9340 39.12 0.9783
-SRFormer-light [80] $\times$2 853K 236G 38.23 0.9613 33.94 0.9209 32.36 0.9019 32.91 0.9353 39.28 0.9785
-Ours $\times$2 859K 198.1G 38.16 0.9610 34.00 0.9212 32.34 0.9017 32.92 0.9356 39.31 0.9779
+LatticeNet [47] $\times$2 756K 169.5G 38.13 0.9610 33.78 0.9193 32.25 0.9005 32.43 0.9302 - -
+SwinIR-light [41] $\times$2 910K 122.2G 38.14 0.9611 33.86 0.9206 32.31 0.9012 32.76 0.9340 39.12 0.9783
+Ours $\times$2 905K 167.1G 38.13 0.9610 33.95 0.9208 32.31 0.9013 32.85 0.9349 39.20 0.9782
 CARN [2] $\times$3 1,592K 118.8G 34.29 0.9255 30.29 0.8407 29.06 0.8034 28.06 0.8493 33.50 0.9440
@@
 LatticeNet [47] $\times$3 765K 76.3G 34.53 0.9281 30.39 0.8424 29.15 0.8059 28.33 0.8538 - -
-SwinIR-light [36] $\times$3 886K 87.2G 34.62 0.9289 30.54 0.8463 29.20 0.8082 28.66 0.8624 33.98 0.9478
-SRFormer-light [80] $\times$3 861K 105G 34.67 0.9296 30.57 0.8469 29.26 0.8099 28.81 0.8655 34.19 0.9489
-Ours $\times$3 867K 88.7G 34.72 0.9296 30.63 0.8475 29.29 0.8099 29.00 0.8689 34.39 0.9495
+SwinIR-light [41] $\times$3 918K 55.4G 34.62 0.9289 30.54 0.8463 29.20 0.8082 28.66 0.8624 33.98 0.9478
+Ours $\times$3 913K 74.5G 34.63 0.9288 30.54 0.8459 29.23 0.8084 28.70 0.8631 34.12 0.9479
 CARN [2] $\times$4 1,592K 90.9G 32.13 0.8937 28.60 0.7806 27.58 0.7349 26.07 0.7837 30.47 0.9084
@@
 LatticeNet [47] $\times$4  [^7] K 43.6G 32.30 0.8962 28.68 0.7830 27.62 0.7367 26.25 0.7873 - -
-SwinIR-light [36] $\times$4 897K 49.6G 32.44 0.8976 28.77 0.7858 27.69 0.7406 26.47 0.7980 30.92 0.9151
-SRFormer-light [80] $\times$4 873K 62.8G 32.51 0.8988 28.82 0.7872 27.73 0.7422 26.67 0.8032 31.17 0.9165
-Ours $\times$4 879K 50.6G 32.51 0.8993 28.85 0.7876 27.75 0.7423 26.75 0.8051 31.26 0.9175
+SwinIR-light [41] $\times$4 930K 31.8G 32.44 0.8976 28.77 0.7858 27.69 0.7406 26.47 0.7980 30.92 0.9151
+Ours $\times$4 924K 42.3G 32.42 0.8977 28.74 0.7847 27.68 0.7400 26.52 0.7983 30.94 0.9135
 Real-world Image Super-resolution.
@@
 The results of gaussian color image denoising are shown in  Tab.
-6.
+5.
 Following [81, 79], the compared noise levels include 15, 25 and 50.
-As one can see, our model achieves better performance than all compared methods.
-In particular, it surpasses the SwinIR [36] by up to 0.25dB with $\sigma$=50 on the Urban100 dataset.
+As one can see, our model achieves the best performance on most datasets.
+In particular, it surpasses the SwinIR [41] by even 0.48dB with $\sigma$=50 on the Urban100 dataset.
 We also give a visual comparison in Fig.
@@
 Restormer [72] 34.40 31.79 28.60 35.47 33.04 30.01 35.61 33.34 30.30 35.13 32.96 30.02
-MambaIR 34.43 31.80 28.61 35.34 32.91 29.85 35.62 33.35 30.31 35.17 32.99 30.07
-MambaIR+ 34.44 31.82 28.64 35.35 32.92 29.87 35.63 33.36 30.32 35.17 32.99 30.06
+MambaIR 34.48 32.24 28.66 35.42 32.99 29.92 35.70 33.43 30.35 35.37 33.21 30.30
 Figure 7: Qualitative comparison of our MambaIR with other methods on color image denoising task with noise level level $\sigma$=50.
@@
 The results, shown in Tab.
-7, suggest that our method achieves comparable performance with existing state-of-the-art models Restormer [61] and outperforms other methods such as Uformer [61] by 0.12dB PSNR on SIDD dataset.
-The promising results indicate the ability of our method in real image denoising.
-Table 7: Quantitative comparison on real image denosing with state-of-the-art methods.
+6, suggest that our method achieves comparable performance with existing state-of-the-art models Restormer [69] and outperforms other methods such as Uformer [69] by 0.12dB PSNR on SIDD dataset, indicating the ability of our method in real image denoising.
+Table 6: Quantitative comparison on the real image denosing task.
 DeamNet [60] MPRNet [73] DAGL [55] Uformer [69] Restormer [72] Ours
@@
 Extensive experiments on multiple restoration tasks demonstrate our MambaIR serves as a simple but effective state-space model for image restoration.
+Acknowledgements
+This work is supported in part by the National Natural Science Foundation of China, under Grant (62302309, 62171248), Shenzhen Science and Technology Program (JCYJ20220818101014030, JCYJ20220818101012025), and the PCNL KEY project (PCL2023AS6-1).